Prove that for all x in R^n and all multi-indices a,
|x^a|<= |x|^(|a|)
1
Expert's answer
2019-10-01T11:06:52-0400

We need to prove that "\frac{|x^a|}{|x|^{|a|}} \leq 1". Rewrite the last expression as:"\frac{|x^{\alpha}|}{|x|^{|\alpha|}} = \frac{ \sqrt{ x_1^{2\alpha_1} + \dots + x_n^{2\alpha_n} } }{|x|^{|\alpha|} } 
\leq ( \frac{ x_1^{2\alpha_1} + \dots + x_n^{2\alpha_n} }{|x|^{2|\alpha|}})^{\frac{1}{2}}
\leq \left(\frac{x_1^{2\alpha_1}}{|x|^{2|\alpha|}} + \dots + \frac{x_n^{2\alpha_n}}{|x|^{2|\alpha|}}\right)^{\frac{1}{2}}
 = \left|\left(\frac{x_1^{\alpha_1}}{|x|^{\alpha_1}}, \dots, \frac{x_n^{\alpha_n}}{|x|^{\alpha_n}}\right )\right|" .

The right side of the last expression is equal to"|e^a|", where "e = (\frac{x_1}{|x|}, ... \frac{x_n}{|x|})" is the unit vector.

Using sup-norm inequality "|x^a| \leq ||x||_{\infty}^{|a|}" and norm inequality "||x||_{\infty}^{|a|} \leq |x|^{|a|}" , obtain "|e^a| \leq |e|^a = 1".

Hence, "\frac{|x^a|}{|x|^{|a|}} \leq 1".